A botanist has developed a new hybrid cotton plant that can withstand insects better than other cotton plants. However, there is some concern about the germination of seeds from the new plant. To estimate the probability that a seed from the new plant will germinate, a random sample of 3000 seeds was planted in warm, moist soil. Of these seeds, 2440 germinated.

(a) Use relative frequencies to estimate the probability that a seed will germinate. What is your estimate? (Enter your answer to 3 decimal places.)


(b) Use relative frequencies to estimate the probability that a seed will not germinate. What is your estimate? (Enter your answer to 3 decimal places.)


(c) Either a seed germinates or it does not. What is the sample space in this problem?

germinate

germinate or not germinate     

not germinate


Do the probabilities assigned to the sample space add up to 1? Should they add up to 1? Explain.

Yes, because they cover the entire sample space.

Yes, because they do not cover the entire sample space.     

No, because they cover the entire sample space.

No, because they do not cover the entire sample space.


(d) Are the outcomes in the sample space of part (c) equally likely?

yes

no    
Answer #1

a)

e probability that a seed will germinate =2440/3000 =0.813

b) probability that a seed will not germinate =1-0.813 =0.187

c) germinate or not germinate     

Yes, because they cover the entire sample space.

d) no